OpenStudy (scott208):

There is a value of a such that subtracting a-4 from 4a+16 gives an answer of -25. What is that value of a?

OpenStudy (welshfella):

form an equation:- 4a + 16 - (a - 4) = -25 Note we put the a - 4 in parentheses because we are subtracting all of this expression.

OpenStudy (welshfella):

solve this for a.

OpenStudy (scott208):

4a+16 -a +4=25

OpenStudy (welshfella):

yes

OpenStudy (scott208):

then 3a+20=25?

OpenStudy (welshfella):

Hold on - i missed that :- right side is -25 NOT 25

OpenStudy (scott208):

ohhhhhhh so 3a=20=-25 and then -20-25=-45/3 and then -15?

OpenStudy (welshfella):

correct

OpenStudy (scott208):

thank ya sir!

OpenStudy (welshfella):

its best to write it out this way - one line for each step:- 3a + 20 = -25 3a = -45 a = -45/3 = -15

OpenStudy (welshfella):

yw